OpenStudy (anonymous):

Write an equation for the line through points (5 , 2)and and ( -2 , 2)

hero (hero):

y = mx + b 2 = 5x + b 2 = -2x + b b = 2 - 5x b = 2 + 2x b = b 2 - 5x = 2 + 2x 2 - 2 = 5x + 2x 0 = 7x x = 0 b = 2 - 5(0) b = 2 y = 2